OpenStudy (anonymous):

If a/b = 4/5, complete this statement: 5/b = ___/___

OpenStudy (anonymous):

@phi

OpenStudy (phi):

if you have an equation like this, you can flip both sides to get \[ \frac{b}{a} = \frac{5}{4} \] that does not help for this question, but remember it. also, you can write a ratio of top (left side) over top (of right side) = bottom (of left) over bottom of right

OpenStudy (anonymous):

So does 5/b = 5/4?

OpenStudy (phi):

so, for example, you can write \[ \frac{a}{4} = \frac{b}{5} \] 5/b = 5/4? NO, you always keep the a,b,4, and 5. just how you write them can change

OpenStudy (anonymous):

I'm still confused though :/

OpenStudy (phi):

when you get time, see this http://www.khanacademy.org/math/arithmetic/rates-and-ratios/ratios_and_proportions/v/introduction-to-ratios--new-hd-version meanwhile, if you have \[ \frac{a}{b}= \frac{4}{5} \] you can also say top left (a) over top right (4) = bottom left (b) over bottom right (5) \[ \frac{a}{4} = \frac{b}{5} \]

OpenStudy (phi):

or we can flip them to get \[ \frac{4}{a} = \frac{5}{b} \]

OpenStudy (anonymous):

So would it be 4/5?

OpenStudy (phi):

look at my last post. 5/b = what ?

OpenStudy (anonymous):

I'm not sure :/ the only thing I can think of is 5/4

OpenStudy (anonymous):

it would be 4/a

OpenStudy (anonymous):

Oh..... well I feel stupid now :/

OpenStudy (phi):

Here is one way to see the ways to write these ratios